SPECTROSCOPY OF SOME N-PHENYL AMIDES. STRUCTURES AND RELATIVE STABILITIES OF THE CIS AND TRANS ISOMERS

\underline{VICTORIA P. MANEA}, JOHN R. CABLE, Department of Chemistry and Center for Photochemical Sciences, Bowling Green State University, Bowling Green, OH 43403.

Equilibrium geometries and isomer distributions in a series of N-phenyl amides were studied using multiphoton ionization spectroscopy in a supersonic jet. The experimental data are consistent with a considerably larger population of the trans isomer in formanilide, although both isomers were present in the expansion. In the cis conformation the phenyl ring is twisted out of the plane of the amide group, while trans-formanilide adopts a planar geometry. Only trans-acetanilide was identified under the same experimental conditions. N-methyl substitution in formanilide results in different cis to trans isomer ratios, with a preference for the cis geometry. In addition, van der Waals clusters of formanilide and acetanilide with water were investigated and evidence of multiple binding sites was seen.